OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that has a slope of 1/4 and contains the point (12,3)

OpenStudy (anonymous):

help needed

OpenStudy (anonymous):

help needed

pooja195 (pooja195):

We need to use the point slope formula \[\huge~y-y_1=m(x-x_1)\] (x1,y1) m=slope

pooja195 (pooja195):

Plug the info into the formula

OpenStudy (anonymous):

9

pooja195 (pooja195):

???

OpenStudy (anonymous):

12-4

OpenStudy (anonymous):

8

OpenStudy (anonymous):

im stuck